# Revisiting Einstein’s analogy: black holes as gradient-index lenses

### v1

@@ VI Results
The photon trajectories in Figs. 57 $(a)$ and $(d)$ were calculated numerically using Eq. (20) in Planck units with $m_{b}=1$, employing the trapezoidal method with 10,000,000 regular divisions and choosing $x_{0}=0.001+0.5r_{bl,+}$. In all cases, the geodesic ends at $R=4$. Planck units were used, and $m_{b}=1$.
@@
In Fig. 5, we present $(a)$ a prograde trajectory ($h=2.8444$) and $(d)$ a retrograde trajectory ($h=-6.832$) for a Kerr black hole with angular momentum $a=0.9$. Based on these trajectories, the analytical refractive index distributions are given by Eq. (30).
@@
where $n(r_{i})$ represents the analytical distribution at point $r_{i}$, $\hat{n}(r_{i})$ is the corresponding numerical reconstruction and $m$ is the total number of discretization points ($i=1,2,...,m$). Lower NRMSE values indicate less residual variance, meaning a better approximation. Here, we choose $m=10,000$.
@@
Another interesting case arises when $a=0$ in Eq. (30) corresponds to the Schwarzschild scenario. In this case, the refractive index distribution is given by:
@@
This result agrees with Ref. 14, although they do not give the explicit form of the constant coefficient.
@@
Finally, a particular case to be considered, as their analytical expression can be obtained straightforwardly, is radial geodesics ($a=h$) described by Eq. (11). These are associated with the following index of refraction:
@@
Figure 7: The radial geodesics ($h=a$): Null geodesics for $(a)$ $h=0.1$ and $(d)$ $h=0.9$, using Planck units and $m_{b}=1$. The corresponding GRIN media are shown in $(b)$ and $(e)$. Figures $(c)$ and $(f)$ compare the exact gradient index (blue line) and PhysGRIN (green dashed line). The GRIN medium is reconstructed between the outer event horizon and $R=4$. The white line indicates the Cauchy horizon.
Figure 8: The gradient index for radial geodesics ($h=a$) for the cases $(a)$ $h=0.1$ and $n_{e}=1/0.0999$, and $(b)$ $h=0.9$ and $n_{e}=1/0.7850$, using Planck units and $m_{b}=1$. The GRIN medium is reconstructed between the outer event horizon and $R=4$. The white line indicates the Cauchy horizon.
All the results presented in this section confirm that GRIN media effectively reproduces the null geodesic and demonstrate the feasibility of using a simple method for reconstructing the GRIN medium without relying on complex techniques. It only requires the ray trajectory through the black hole. This highlights the simplicity of the exact numerical method, which is one of the main objectives of the present work.
VII Unstable GRIN media for radial geodesics of the Schwarzschild black hole
Numerically, issues are expected when computing the GRIN medium for trajectories close to the radial geodesic ($h\rightarrow 0$), as in this case $K\rightarrow 0$, which may introduce errors in the numerator of Eq. (35) under such conditions.
On the other hand, from the analytical expression in Eq. (39), when $h=0$ and $r\neq 0$, the index of refraction remains constant and equal to $n_{e}$, which is expected for a straight light ray moving toward the origin, interestingly, in Ref. 14, although their expression for the index of refraction is nearly identical, for the radial geodesic it tends to infinity throughout the space. This discrepancy arises because their unspecified constant of proportionality does not seem to depend on $h$.
@@ VII Unstable GRIN media for radial geodesics of the Schwarzschild black hole
This results in yet another different expression for the index of refraction for the radial geodesic. In this case, the index of refraction is well-defined at $r=0$, which is inconsistent with the black hole model, as it allows light to traverse the black hole.
@@
These discrepancies in the index of refraction when $h=0$ and $a=0$ demonstrate that there is no well-defined limit for the index of refraction for a radial geodesic in the Schwarzschild black hole, indicating that the index of refraction is unstable for radial geodesics. In experiments, this instability means that null geodesics close to the radial geodesics should be avoided.
VIII Discussion and Conclusions
@@
In this work, we presented a methodology based on the numerical PhysGRIN method to derive GRIN media replicating photon trajectories in the curved spacetime around equatorial Kerr black holes. This approach leverages the system’s symmetry and known photon trajectories, enabling the construction of GRIN media without requiring advanced knowledge of general relativity or tensor calculus. The method is founded on the conservation of Fermat’s ray invariant, offering a robust and flexible numerical tool for modeling complex optical media.
We used Cartesian coordinates to represent ray trajectories because they represent the Kerr black hole singularity as an annulus, a widely accepted depiction.
@@
Additionally, by applying the conservation of Fermat’s ray invariant, we derived an analytical form of the GRIN medium for equatorial null geodesics around Kerr black holes in Cartesian coordinates. In contrast, Ref. 14 provided expressions in Boyer-Lindquist coordinates, differing from ours except in the Schwarzschild case, where they coincide up to a constant, leading to additional issues for radial geodesics.
@@
Our results align with the suggestion of Ref. 18 that the most accurate definition of the index of refraction compares null geodesics in spacetime to light rays in optical media. This approach is particularly effective when focusing solely on ray trajectories without considering additional electromagnetic wave properties such as polarization.
Overall, the PhysGRIN method provides a framework for deriving GRIN media from known light paths, producing an index of refraction that matches the system’s geodesics. While the results depend on how light rays are represented in a specific coordinate system, this approach remains independent of alternative conceptual interpretations of the index of refraction, as evidenced by the various definitions proposed for the Schwarzschild case Ramezani_Aval_2024 . By relying solely on the symmetries of light paths, the method can also be extended to mimic particle trajectories in classical and quantum mechanics (including Bohmian trajectories) Xiao_2021 ; de_Alcantara_Bonfim_2000 . This makes GRIN media a valuable tool for visualizing phenomena across various scales, from quantum systems to macroscopic scenarios.
Disclosures
The authors declare that there are no financial interests, commercial affiliations, or other potential conflicts of interest that could have influenced the objectivity of this research or the writing of this paper.
Code Availability
Data may be obtained from the authors upon reasonable request.
Acknowledgments
The authors acknowledge the support of the Instituto Nacional de Astrofísica, Óptica y Electrónica (INAOE), the Consejo Nacional de Humanidades, Ciencias y Tecnologías (CONAHCYT), and the Sistema Nacional de Investigadoras e Investigadores (SNII) for this publication.

### v2

@@ VI Results
The photon trajectories in Figs. 57 $(a)$ and $(d)$ were calculated numerically using Eq. (20) in Planck units with $m_{b}=1$, employing the trapezoidal method with 10,000,000 regular divisions and choosing $x_{0}=0.001+0.5r_{bl,+}$. In all cases, the geodesic ends at $R=4$.
@@
[…]
@@
where $n(r_{i})$ represents the analytical distribution at the point $r_{i}$, $\hat{n}(r_{i})$ is the corresponding numerical reconstruction and $m$ is the total number of discretization points ($i=1,2,...,m$). Lower NRMSE values indicate less residual variance, meaning a better approximation. Here, we choose $m=10,000$.
@@
Another interesting case arises when $a=0$ in Eq. (30), which corresponds to the Schwarzschild scenario. In this case, the refractive index distribution is given by:
@@
This result agrees with Ref. Tinguely_2020 , although they do not give the explicit form of the constant coefficient.
@@
Finally, a particular case to be considered, since its analytical expression can be obtained straightforwardly, is when the geodesics have a vanishing Carter constant ($a=h$) described by Eq. (11). These are associated with the following index of refraction:
@@
Figure 7: Geodesics with a vanishing Carter constant ($h=a$): Null geodesics for $(a)$ $h=0.1$ and $(d)$ $h=0.9$, using Planck units and $m_{b}=1$. The corresponding GRIN media are shown in $(b)$ and $(e)$. Figures $(c)$ and $(f)$ compare the exact gradient index (blue line) and PhysGRIN (green dashed line). The GRIN medium is reconstructed between the outer event horizon and $R=4$. The white line indicates the Cauchy horizon.
Figure 8: The gradient index for geodesics with a vanishing Carter constant ($h=a$) for the cases $(a)$ $h=0.1$ and $n_{e}=1/0.0999$, and $(b)$ $h=0.9$ and $n_{e}=1/0.7850$, using Planck units and $m_{b}=1$. The GRIN medium is reconstructed between the outer event horizon and $R=4$. The white line indicates the Cauchy horizon.
All the results presented in this section confirm that GRIN media effectively reproduce the null geodesic and demonstrate the feasibility of using a simple method for reconstructing the GRIN medium without relying on complex techniques. It only requires the ray trajectory through the black hole spacetime. This highlights the simplicity of the exact numerical method, which is one of the main objectives of the present work.
VII Unstable GRIN media for radial null geodesics of the Schwarzschild black hole
In the Schwarzschild metric ($a=0$), a radial null geodesic is a straight line going to the center of the black hole $(\theta=\text{constant}$, $\phi=\text{constant}$). When the rays are considered to lie in the equatorial plane $(\theta=\pi/2$) with the source at $r=R$, $\phi=0$, numerical issues are expected when computing the GRIN medium for trajectories close to the radial geodesic (i.e., $h\to 0$), as in this case, $K_{r}\to 0$ in Eq. (33), which may introduce errors in the numerator of Eq. (35) under such conditions.
On the other hand, from the analytical expression in Eq. (39), when $h=0$ and $r\neq 0$, the index of refraction remains constant and equal to $n_{e}$, which is expected for a straight light ray moving toward the origin, interestingly, in Ref. Tinguely_2020 , although their expression for the index of refraction is nearly identical, for the radial geodesic it tends to infinity throughout the space. This discrepancy arises because their unspecified constant of proportionality does not seem to depend on $h$.
@@ VII Unstable GRIN media for radial null geodesics of the Schwarzschild black hole
This results in yet another different expression for the index of refraction for the radial geodesic. In this case, the index of refraction is well-defined at $r=0$, which is inconsistent with the black hole model as it allows light to traverse the black hole.
@@
These discrepancies in the index of refraction when $h=0$ and $a=0$ demonstrate that there is no well-defined limit for the index of refraction for a radial geodesic in the Schwarzschild spacetime. This indicates that the index of refraction is unstable for radial geodesics. In experiments, this instability means that null geodesics close to the radial geodesics should be avoided.
VIII Robustness of PhysGRIN under stochastic perturbations
To evaluate the robustness of the PhysGRIN method, a null geodesic with vanishing Carter constant and parameters $h=a=0.9$ is considered as a representative example. The geodesic starts at radius $x_{0}=0.001+0.5r_{bl,+}$, extends up to a radial distance of $R=4$, and is discretized into $p$ data points. Each point is randomly displaced along the normal direction, introducing a standard deviation $\sigma$.
The PhysGRIN method relies on Eq. (35), which depends directly on the sine of the angle $\theta_{Inc}$ of the incident straight light ray over the GRIN surface with respect to its normal, and inversely on $\sin(\varphi_{i})$, where $\varphi_{i}$ is defined in Eq. (36). Because of this, directly connecting the displaced points to form a trajectory and then applying the PhysGRIN method can lead to problems as shown in Fig 9. Figure 9$(a)$ shows the randomly displaced points (yellow dots) for $\sigma=0.1$ and $p=100$. The corresponding GRIN medium in Fig. 9$(b)$ yields an NRMSE of 1.1175 relative to the unperturbed case, indicating low reconstruction accuracy. Reducing the noise to $\sigma=0.01$ while keeping $p=100$ (Fig. 9$(c)$) significantly improves the result, with an NRMSE of 0.0667 (Fig. 9(d)). However, increasing the number of points to $p=1000$ while keeping $\sigma=0.01$ (Fig. 9$(e)$) leads to a significant drop in accuracy, with an NRMSE of 2.9081 (Fig. 9$(f)$).
Figure 9: The exact null geodesic (green curve), with vanishing Carter constant and $a=h=0.9$, is shown along with $p$ stochastic points (yellow) distributed with standard deviation $\sigma$ from the geodesic: $(a)$ $\sigma=0.1$, $p=100$; $(c)$ $\sigma=0.01$, $p=100$; and $(e)$ $\sigma=0.01$, $p=1000$. The corresponding reconstructed gradient index profiles are shown on the right (yellow curve and colormap) in figures $(b)$, $(d)$, and $(f)$, along with the exact gradient index (blue curve). Circular red dashed lines mark the exterior boundary of the GRIN media. Red dashed lines indicate the radial region where the exact null geodesic lies. Planck units are used, and $m_{b}=1$.
When the angles $\varphi_{i}$ change smoothly between points, as seen in the zoomed view of Fig. 9(c), the index of refraction is estimated more accurately (Fig. 9(d)). However, abrupt changes in consecutive values of $\varphi_{i}$ cause large variations in the index of refraction. In extreme cases, when $\varphi_{i}$ approaches zero or $\pi$ (Fig. 9(e)), the index can become very large or even diverge (Fig. 9(f)). Moreover, $\theta_{Inc}$ is linked with the initial direction of the ray geodesic, and at this point, we know that for every trajectory we obtain a different index of refraction; thus, taking a random initial value of $\theta_{Inc}$ also leads to error.
Figure 10: The fitted geodesic (purple curve), obtained via least-squares minimization from $p$ stochastic points (yellow) with standard deviation $\sigma$ from the exact geodesic (with vanishing Carter constant and $a=0.9$), is shown for: $(a)$ $\sigma=0.1$, $p=100$; $(c)$ $\sigma=0.1$, $p=1000$; and $(e)$ $\sigma=0.01$, $p=1000$. Figures $(b)$, $(d)$, and $(f)$ on the right show the corresponding reconstructed gradient index profiles (dashed purple curve and colormap), along with the exact gradient index profile (blue curve). Red dashed lines indicate the GRIN media boundaries. Planck units are used, and $m_{b}=1$.
Taking control of this issue, or when dealing with stochastic data such as experimental trajectories of stars or planets (since PhysGRIN is directly applicable to massive particles, as will be shown in Section IX), a suitable approach is to first fit a smooth curve to the data. This helps avoid abrupt changes in the index of refraction, which are not physically expected.
To build a smooth curve from the stochastic points, each data point $p$ is represented in polar coordinates as $(\theta_{\text{s},p},r_{\text{s},p})$. The data suggest an exponential dependence on the angular coordinate $\theta_{\text{s},p}$, so the points are reparametrized as $(\exp(\theta_{\text{s},p}),r_{\text{s},p})$. The curve is then fitted using a fifth-degree polynomial via least-squares minimization burden2011numerical .
The left column of Fig. 10 shows the reconstructed ray (purple curve) from stochastic data for three cases: $\sigma=0.1$ with $p=100$ (Fig. 10$(a)$), $\sigma=0.1$ with $p=1000$ (Fig. 10$(c)$), and $\sigma=0.01$ with $p=1000$ (Fig. 10$(e)$). The corresponding GRIN media are shown on the right (Figs. 10$(b)$, $(d)$, and $(f)$). The NRMSE values are $0.0892$, $0.0844$, and $0.0056$, respectively. This clearly demonstrates a significant improvement, and therefore shows that the PhysGRIN method is robust against stochastic perturbations, provided that the trajectory is pre-smoothed (using, for instance, least-squares fitting), thereby enabling reliable and accurate reconstruction of complex GRIN media.
IX Reconstructing the S2 Orbit via PhysGRIN: From Newtonian to Relativistic Models
Up to this point, the capability of the PhysGRIN method to reconstruct the index of refraction profile that reproduces photon trajectories in the equatorial plane of a Kerr black hole has been demonstrated, including the particular case of the Schwarzschild spacetime. The method requires only the trajectory itself, making it independent of the technique used to obtain the path and, therefore, also independent of the nature of the particle; i.e., it can be applied to any object whose path through spacetime can be tracked.
In this section, the capability and accuracy of the PhysGRIN method in distinguishing subtle relativistic deviations from Newtonian predictions in the motion of massive particles are evaluated through the analysis of the apsidal precession of the star S2, which orbits the nearest supermassive black hole candidate at the center of the Milky Way 2020 . To distinguish this relativistic effect from the Newtonian gravitational potential case, three orbital models are considered: (i) a Newtonian elliptical orbit, (ii) a quasi-elliptical trajectory that includes precession, and (iii) the relativistic orbit obtained via numerical integration.
Figure 11: In figure $(a)$, half of the trajectory followed by the S2 star without precession is shown (green curve), along with the index of refraction profile obtained by applying the PhysGRIN method. Figure $(b)$ shows the exact index of refraction profile (blue curve) and the reconstructed profile (dashed green curve). Here, lengths are expressed in astronomical units.
(i) The Newtonian elliptic orbit. In short, within classical mechanics, the conserved energy per unit mass, $E_{N}$, and the orbital angular momentum per unit mass, $L_{N}$, for motion in the equatorial plane of a central potential satisfy goldstein2002classical :
$2E_{N}=\left(\frac{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\text{d}}r}{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\text{d}}t}\right)^{2}+\frac{L_{N}^{2}}{r^{2}}+2V(r),\quad L_{N}=r\frac{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\text{d}}\phi}{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\text{d}}t},$ (43)
where, for an elliptical orbit in the Newtonian potential $V(r)=-GM/r$, the specific energy and angular momentum are goldstein2002classical
$E_{N}=-\frac{GM}{2a_{orb}}\quad\text{and}\quad L_{N}=\sqrt{GMa_{orb}(1-e_{orb}^{2})},$ (44)
with $a_{orb}$ denoting the semi-major axis of the orbit and $e_{orb}$ its eccentricity. The position of the star S2 as a function of its angular position in the equatorial plane is given by the standard conic section equationgoldstein2002classical
$r(\phi)=\frac{a_{orb}(1-e_{orb}^{2})}{1+e_{orb}\cos\phi},$ (45)
Due to orbital symmetry, only half of the trajectory is sufficient to reconstruct the gradient index distribution using the PhysGRIN method, which reproduces the star’s path. By taking $M=4.2997\times 10^{6},M_{\odot}$, $a=1034.6$ AU, and $e=0.88444$ Ashkenazy_2025 , the PhysGRIN method achieves a normalized root-mean-square error (NRMSE) of $7.3579\times 10^{-7}$ with respect to the analytical index of refraction profile obtained through the optical-mechanical analogy leonhardt2010geometry ; Xiao_2021 . This analytical profile is computed using the Newtonian potential and normalized such that the index of refraction equals $1$ at the apoapsis. Figure 11 shows the reconstructed gradient index profile obtained with PhysGRIN, along with the exact index of refraction, revealing an almost perfect match.
(ii) Quasi-elliptical orbit accounting for precession. As a first approach to modeling an orbit with apsidal precession, a representation of the trajectory is obtained through a quasi-elliptical path by replacing the term $\cos\phi$ in Eq. (45) with $\cos(\sqrt{K_{p}}\phi)$, following Hall_2022 ; rindler2006relativity , where
$K_{p}=1-\frac{6GM}{c^{2}a_{orb}(1-e_{orb}^{2})}.$ (46)
Although this approximation is valid for orbits with low eccentricity, which deviates only slightly from a circular orbit. Such a modification yields an apsidal shift of
$2\pi\left(\frac{1}{\sqrt{K_{p}}}-1\right)\approx 12.2^{\prime}.$ (47)
a value that does not differ significantly from the $12.1^{\prime}$ reported in 2020 . Such a precession effect is not easy to detect over a single orbital period, making it difficult to distinguish from an elliptical trajectory. In the context of an optical device with dimensions on the order of meters, reproducing precession would be particularly challenging. Although more precise values for the precession can be obtained by accounting for the stellar distribution around SgrA* Ashkenazy_2025 , the objective here is to assess whether sufficient accuracy can be achieved via the PhysGRIN method to differentiate between slightly different trajectories. Therefore, the present model is adequate for such analysis.
The PhysGRIN reconstruction for this quasi-elliptic orbit yields an NRMSE of $3.3219\times 10^{-4}$ with respect to the exact index of refraction reported in Xiao_2021 for the modified Newtonian potential $V(r)=-GMr^{-1}-GML^{2}c^{-2}r^{-3}$ rindler2006relativity , thus exhibiting greater error than in the previous case of the Newtonian elliptic orbit (on the order of $10^{-7}$). This increased error can be attributed to the quasi-elliptic approximation used for modeling the orbit of S2.
(iii) Relativistic orbit obtained via numerical integration. From Eq. (43), the angular position as a function of radius is given by
$\phi(r)=\int_{r_{I}}^{r_{f}}\frac{L\,{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\text{d}}r}{r^{2}\sqrt{2E-L^{2}r^{-2}-2V(r)}}$ (48)
To account for the relativistic effect of apsidal precession, it is sufficient to use the modified Newtonian potential in the previous integral. To avoid numerical issues when using the trapezoidal method for integration (due to complex values arising inside the square root) the integration limits are chosen as $r_{I}=1949.6$ and $r_{f}=119.5$, values close to the apoapsis and periapsis distances of the S2 star relative to SgrA*, respectively. The NRMSE associated with this numerical trajectory, compared to the exact gradient index for the modified Newtonian potential Xiao_2021 , is $2.3133\times 10^{-6}$, thus showing less error than the quasi-elliptic precession orbit.
Figure 12: Subtraction of refractive index profiles between the elliptical trajectory and those including apsidal precession for: (a) exact analytical indices from the optical-mechanical analogy leonhardt2010geometry ; Xiao_2021 (blue line); (b) PhysGRIN reconstructions of the elliptical and quasi-elliptical trajectories (yellow line); and (c) PhysGRIN reconstructions of the elliptical and numerically computed relativistic trajectories (green dashed line). In all cases, 1,000,000 points were used to generate the corresponding orbit trajectory. Here, length is expressed in astronomical units.
Figure 12 shows the subtraction between the refractive index profiles corresponding to the elliptical trajectory and those including apsidal precession, for the following cases: (a) the subtraction of the exact analytical indices of refraction obtained via the optical-mechanical analogy leonhardt2010geometry ; Xiao_2021 between the Newtonian and modified Newtonian potentials (blue line); (b) the subtraction of the PhysGRIN reconstruction between the elliptical trajectory (case (i)) and the quasi-elliptical trajectory (case (ii)), shown by the yellow line; and (c) the subtraction of the PhysGRIN reconstructions between the elliptical trajectory (case (i)) and the numerically computed relativistic trajectory of the S2 star (case (iii)), shown by the green dashed line. Although all models yield low NRMSE values, the differences between curves (b) and (c) in Fig. 12 are clearly noticeable. Resolving such variations requires reconstruction errors below $10^{-5}$. Therefore, in practice, fabricating a GRIN medium capable of accurately reproducing the apsidal precession of the S2 star would be challenging.
X Conclusions
@@
In this work, we presented a methodology based on the numerical PhysGRIN method to derive GRIN media replicating photon trajectories in the equatorial plane of the Kerr spacetime. This approach leverages the system’s symmetry and known photon trajectories, enabling the construction of GRIN media without requiring advanced knowledge of GR or tensor calculus while remaining independent of the method used to calculate the null geodesics (whether analytic, numeric, or approximate).
While PhysGRIN relies on the conservation of Fermat’s ray invariant via Noether’s theorem (providing a robust and flexible numerical tool for modeling complex optical media), the theory presented here can be extended to model light rays outside the equatorial plane of the Kerr spacetime and potentially for more exotic metrics, using a three-dimensional freeform GRIN medium Lippman_2021 . Due to the complexity of this task, this topic will be addressed in a future paper. Note that although photon trajectories are not directly observable, gravitational lensing can give us indirect experimental verification of such theoretical trajectories Lemon_2024 ; furthermore, we can also account for the deflection phenomena of massive particles due to the spacetime curvature using GRIN media, as shown in Xiao_2021 , or by applying the method presented in this work. Moreover, relativistic particles, such as neutrinos emitted by the accretion disk of a massive compact object, could help detect their shadows, as suggested in Mendoza_2024 . These trajectories can also be modeled using GRIN media with the approach presented in this paper.
@@
Additionally, by applying the conservation of Fermat’s ray invariant, we derived an analytical form of the GRIN medium for equatorial null geodesics around Kerr black holes in Cartesian coordinates (we used Cartesian coordinates to represent ray trajectories as they represent the Kerr black hole singularity as an annulus, a widely accepted depiction). In contrast, Ref. Tinguely_2020 provided expressions in Boyer-Lindquist coordinates, differing from ours except in the Schwarzschild case, where they coincide to a constant, leading to additional issues for radial geodesics. Moreover, the approach presented here goes beyond the limitation of mapping the spatial part of the metric to a conformally flat coordinate system to obtain the index of refraction that exactly reproduces photon trajectories, as suggested in Parvizi_2024 . Thus, PhysGRIN offers a more general approach to reconstruct GRIN media.
@@
Our results align with the suggestion of Ref. Ramezani_Aval_2024 that the most accurate definition of the index of refraction compares null geodesics in spacetime to light rays in optical media. This approach is effective when concentrating solely on ray trajectories without considering additional electromagnetic wave properties such as polarization.
An extension of this work would be the application of PhysGRIN to model light ray propagation around black holes surrounded by plasma. Plasma near a black hole alters the equations of motion for photons. One way to account for such deviations in photon trajectories is by introducing refractive properties of the spacetime through an index of refraction that depends on the position and the frequency of the light Atamurotov_2015 ; Rogers_2015 . If the plasma distribution confines light rays to the equatorial plane, the method introduced here to reconstruct GRIN media with spherical symmetry is applicable, but now in a non-normalized GRIN medium Gomez-Correa:22 (a normalized GRIN medium is understood when the refractive index at its surface matches that of the surrounding medium). On the other hand, when plasma is present, additional effects arise due to light-matter interactions. While we can conjecture that light-matter interactions could be modeled by incorporating dispersion into the GRIN medium, it is important to note that light polarization (such as those arising from the presence of an accretion disk Matt_1993 ; Beloborodov_1998 ) is irrelevant in the context of light propagation in a two-dimensional plane, which is the case of light propagating in the equatorial plane of a Kerr black hole. However, when considering more general photon trajectories, a bianisotropic medium that reproduces null geodesics and accounts for polarization must be used Tinguely_2020 . Optical analogs of such astrophysical systems require further study, which is beyond the scope of this manuscript.
Another phenomenon that could be modeled using GRIN media is the propagation of gravitational waves. In the high-frequency regime—when the wavelength is much smaller than the characteristic curvature scale of spacetime—gravitational waves propagate along null geodesics, analogous to light rays in geometrical optics. This approximation, developed by Isaacson in 1968 through a WKB-type expansion of Einstein’s field equations, remains valid even in strong gravitational fields and describes gravitational waves as spin-2 perturbations traveling on a curved background Isaacson1968 . Under these conditions, light treated as a scalar wave in a GRIN medium can effectively emulate their behavior.
Linking ray optics and wave optics through integral diffraction catastrophes is a common practice Berry_1980 ; this, however, must be treated with caution in the case of gravitational lensing for the following reasons: First, observing interference phenomena in real gravitational lensing scenarios is challenging due to coherence issues arising from the finite size of sources or the lack of near-perfect alignment, as mentioned in Rodr_guez_Fajardo_2023 . On the other hand, while gravitational waves might be better suited for observing diffraction and interference phenomena experimentally, they are only pertinent for gravitational waves with wavelengths comparable to the Schwarzschild radius; however, this is beyond geometric optics Cusin_2020 . Second, the GRIN media approach introduced in this paper depends on the trajectory of the photon; hence, only narrow, highly directional beams (such as Gaussian beams) are expected to best approximate geodesic trajectories, as shown in Tinguely_2020 . This means that wave phenomena can only be analyzed within a narrow region. Addressing the connection between ray optics and wave optics requires optical metamaterials or freeform inhomogeneous media capable of capturing the global behavior of photon trajectories.
To conclude, the PhysGRIN method introduced in this work provides a general and systematic approach to constructing GRIN media that accurately reproduce null geodesics in curved spacetimes. By relying on the symmetries of light paths, the method enables the design of refractive index distributions without the need to map the spacetime metric to a conformally flat form, thus offering a versatile tool for modeling light propagation in both relativistic and analog optical systems. Moreover, its trajectory-based formulation allows the method to be extended to simulate particle dynamics in other physical contexts, including classical mechanics and quantum systems such as Bohmian trajectories. When dealing with stochastic data, the method remains robust by fitting a smooth curve that mitigates errors from spurious values in the expected refractive index caused by abrupt changes in ray direction. The approach is simple, practical, and sufficiently precise to distinguish subtle differences between trajectories, as shown in the comparison of orbital models with and without the apsidal precession of the S2 star around SgrA*.
Acknowledgements.
The corresponding author thanks the Secretaría de Ciencia, Humanidades, Tecnología e Innovación (SECIHTI) for the support provided through the awarded postdoctoral scholarship (CVU 640385).